OpenStudy (anonymous):

how do you solve x+4/7=5

OpenStudy (anonymous):

uhmm make the denominator be the same for all so it becomes 7x + 4 = 35

OpenStudy (anonymous):

then 7x = 35 - 4 so x = 31/7

OpenStudy (anonymous):

thanks man

OpenStudy (anonymous):

=]

OpenStudy (anonymous):

another question how do you do 2x/3+8=20

OpenStudy (anonymous):

same thing, make the denominator same through out meaning x3 all so u get 2x + 24 = 60 so 2x = 60 -24 so finally x = 36/2 x = 18

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

(2x/3)*3+8*3=20*3 2x + 24 = 60 x = 36/2 x = 18

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

how do you do 11x+7=3x-9

OpenStudy (stacey):

Get the x's together on one side of the equal sign. It can be done by subtracting 3x from both sides. Then solve as was done with the other problems.